Person A travels 2400 miles in 4 hours 16 minutes. Person B travels the same distance at a speed 2/3 as fast as Person A. What is the speed of Person B
Answer by Alan3354(69443)

B would travel 2/3 of 2400 miles in 4:16 then.
1600 mi/(4 4/15 hrs) = 1600/(64/15)
= 1600*15/64
= 375 mi/hr